OpenStudy (anonymous):

A piece of wire x inches long will be bent into the shapes below. Write an expression for the enclosed area of each shape in terms of x. a) Circle: A=_______ square inches. b) Square: A=_______ square inches.

OpenStudy (jhannybean):

Here you go. http://www.mathsisfun.com/area.html

OpenStudy (anonymous):

@Jhannybean I'm still struggling with this.

OpenStudy (anonymous):

I tried setting x equal to the formula of the circumference and then solving for r and substituting into the area formula.

OpenStudy (jhannybean):

One minute D:

hero (hero):

@iamsammybear...what do you mean 'you tried'? Can you show your work please?

OpenStudy (anonymous):

\[x=2 \pi r^{2}\]\[r=\sqrt{x-2\pi}\]\[A=\pi (x-2\pi)\]

hero (hero):

you made a mistake while solving for r.

hero (hero):

Your formula for circumference is also incorrect

hero (hero):

\(x = C = 2\pi r\) \(x = 2\pi r\) \[r = \frac{x}{2\pi}\]

OpenStudy (jhannybean):

Alright. in order to find the Area of the circle in terms of x, first we need to fidn the circumference....since we're given a wire and we want to see HOW MUCH of the wire it would take to circumscribe the circle. C = \(2 \pi r\) Putting this in terns of x, we can say \(x= 2 \pi r\) therefore \(r= x/2\pi\) Now that we've found our radius, we can plug it into the equation of a circle for Area. \[\large A = \pi r^2\]\[\large A = \pi \left( \frac{x}{2\pi}\right)\]\[\large A = \frac{x}{2}\]

OpenStudy (anonymous):

Oh, I see. Thank you so much @Hero!

OpenStudy (jhannybean):

oh no! i forgot the squared.

hero (hero):

@Jhannybean, you made a mistake in your simplifcation

OpenStudy (jhannybean):

Yes i know..... \[\large A = \pi \left(\frac{x}{2\pi}\right)^2\]\[\large A = \pi \left(\frac{x^2}{4\pi^2}\right)\]\[\large A = \frac{x^2}{4\pi}\]

OpenStudy (anonymous):

Thank you all so much! I really appreciate the help from all of you.

OpenStudy (jhannybean):

And then for the square.. \[\large A = x = 4s \]\[\large s = \frac{x}{4}\] Now plug it into the formula for Area of a square, A = s^2 \[\large A = \left(\frac{x}{4}\right)^2 = \frac{x^2}{16}\]

hero (hero):

@Jhannybean, (as a suggestion) you should have allowed @iamsammybear to try the other one on her own.

OpenStudy (jhannybean):

You are right....